# Calibrated Probabilistic Interpolation for GEDI Biomass

### v2

@@ 4 Discussion
A central contribution of this work is demonstrating that ensemble-based methods produce miscalibrated uncertainty estimates for biomass mapping that this miscalibration stems from two distinct but related issues: (1) a statistical distinction between ensemble variance and prediction intervals, and (2) an inability to adapt uncertainty to local spatial context in heterogeneous landscapes. We discuss each in turn.
@@ 4.2 Context-Aware Uncertainty for Heterogeneous Landscapes
XGBoost in our implementation has an alternative method of uncertainty estimation through quantile regression. Instead of predicting the mean, the model learns quantile functions by minimizing pinball loss. This approach directly targets prediction intervals rather than relying on ensemble variance. The algorithm builds trees sequentially, with each new tree fitting the residuals of the ensemble so far:
@@ 5.4 Propagating Observation-Level Uncertainty
This extension would enable the model to distinguish between predictions that are uncertain because nearby GEDI shots are unreliable versus predictions that are uncertain because the landscape is heterogeneous. Such disentanglement could improve interpretability for operational users and provide more principled uncertainty estimates when GEDI sampling quality varies across a study area. However, implementation requires careful consideration of how GEDI uncertainties are themselves derived and calibrated, as propagating miscalibrated input uncertainties could compound rather than correct biases in the final predictions.
@@ Appendix B Implementation Details
Using the gedidb library, we filter shots based on beam sensitivity, surface detection quality, and agreement with TanDEM-X elevation data. The final training dataset is constructed via an inner join on shot identifiers. This removes observations in dense canopies where ground detection may be unreliable, even if the L4A product does not explicitly flag them.

### v3

@@ 4 Discussion
A central contribution of this work is demonstrating that ensemble-based methods produce miscalibrated uncertainty estimates for biomass mapping that this miscalibration stems from two distinct but related issues: (1) a statistical distinction between ensemble variance and prediction intervals, and (2) an inability to reflect local spatial context in heterogeneous landscapes in uncertainty estimates. We discuss each in turn.
@@ 4.2 Context-Aware Uncertainty for Heterogeneous Landscapes
XGBoost, in our implementation, has an alternative method of uncertainty estimation through quantile regression. Instead of predicting the mean, the model learns quantile functions by minimizing pinball loss. This approach directly targets prediction intervals rather than relying on ensemble variance. The algorithm builds trees sequentially, with each new tree fitting the residuals of the ensemble so far:
@@ 5.4 Propagating Observation-Level Uncertainty
This extension would enable the model to distinguish between predictions that are uncertain because nearby GEDI shots are unreliable versus predictions that are uncertain because the landscape is heterogeneous. Such disentanglement could improve interpretability for operational users and provide more principled uncertainty estimates when GEDI sampling quality varies across a study area. However, implementation requires careful consideration of how GEDI uncertainties are themselves derived and calibrated, as propagating miscalibrated input uncertainties could compound rather than correct biases in the final predictions. We plan to explore this in future work.
@@ Appendix B Implementation Details
Using the gedidb library (Besnard et al., 2025), we filter shots based on beam sensitivity, surface detection quality, and agreement with TanDEM-X elevation data. The final training dataset is constructed via an inner join on shot identifiers. This removes observations in dense canopies where ground detection may be unreliable, even if the L4A product does not explicitly flag them.